# Joint Modelling of Astrophysical Systematics for Cosmology with LSST

### v2

@@ 2.3 Intrinsic alignment of galaxies
The parametrisation of the redshift and luminosity dependent IA amplitude used in this work builds upon Joachimi et al. 2011 and Krause et al. 2015:
$A(L,z)=A_{0}\frac{C_{1}\rho_{m,0}}{D(z)}\left(\frac{L}{L_{0}}\right)^{\beta}\left(\frac{1+z}{1+z_{0}}\right)^{\eta},$ (8)
where $A_{0}$ is a prefactor and $z_{0}$ is a pivot redshift. $C_{1}$ is a normalisation constant (fixed to $C_{1}\rho_{\mathrm{cr}}\approx 0.0134$, with $\rho_{\mathrm{cr}}$ the critical density), $\rho_{\mathrm{m},0}$ is the matter density at $z=0$ (today), and $D(z)$ is the linear growth factor. We follow Eq. 8 from Krause et al. 2015 where the amplitude is averaged over the population of red galaxies:
$A(m_{\mathrm{lim}},z)=\biggl{\langle}A(L,z)\biggr{\rangle}_{\mathrm{red}}\times\left[\Theta(\tilde{z}-z)+\Theta(z-\tilde{z})\left(\frac{1+z}{1+\tilde{z}}\right)^{\eta_{z}^{\mathrm{high}}}\right]$ (9)
with $m_{\mathrm{lim}}$ the survey limiting magnitude, $\Theta$ the step function where the redshift scaling goes from low to high regime at some pivotal redshift $\tilde{z}$ and with
@@
$A(m_{\mathrm{lim}},z)=A(L_{0},z)\left\langle\left(\frac{L}{L_{0}}\right)^{\beta}\right\rangle_{\phi_{\mathrm{red}}}=$
$A(L_{0},z)\frac{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\left(\frac{L}{L_{0}}\right)^{\beta}\phi_{\mathrm{red}}(L,z)}{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\phi_{\mathrm{red}}(L,z)}.$ (10)
The fraction of red galaxies is obtained via:
$f_{\mathrm{red}}=\frac{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\phi_{\mathrm{red}}(L,z)}{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\phi_{\mathrm{all}}(L,z)}$ (11)
with $\phi_{\mathrm{red}}$ and $\phi_{\mathrm{all}}$ the luminosity functions of for red and all galaxies in the sample.
@@
An important emphasis is that the IA amplitude is modeled as luminosity-dependent. Thus, when computing the cosmic shear spectrum from theory, it is necessary to assume a luminosity function parameterisation. Previous forecasting work using this parameterisation has fixed the luminosity function parameters to best-fit values measured from external data sets (Krause et al., 2015; Mandelbaum et al., 2018). However, it is not clear that this is an appropriate analysis choice; we will investigate this as part of the development of our analysis framework below in Sec. 3.
@@ 2.4 Cosmic shear
Each contribution is described in spherical harmonic space where $\ell$ is the angular frequency (a 2D wave vector perpendicular to the line of sight) while indices $(i,j)$ represent tomographic redshift bins where the shear correlations are computed (for a review of key intrinsic alignment quantities and notation, see Lamman et al. 2024).
In this work, we will take our fiducial IA model to be the non-linear alignment (NLA) model (Bridle & King, 2007; Krause et al., 2015). We stress that although it is likely that a more complex model of IA may be required for LSST analysis (e.g. the TATT model from Blazek et al. 2019), the NLA model provides a suitable means of incorporating luminosity-dependent IA modelling and is appropriate in the context of a first exploration of our joint modelling methodology. For the case of the non-linear alignment model of intrinsic alignments, each term in the observed cosmic shear power spectrum is given by:
@@
where in the linear alignment model, $P_{\epsilon\epsilon}$ and $P_{\delta\epsilon}$ are given by the linear matter power spectrum, whilst the non-linear alignment model uses the non-linear matter power spectrum. Both linear and non-linear matter power spectra are expressed in the Fourier space, $k$. $A(z)$ is the redshift dependent intrinsic alignment amplitude that also depends on the galaxy properties and we use the amplitude parametrisation presented in Sec. 2.3.
@@ 2.5 Joint modelling via the luminosity function
Given the introduction of the convolutional method of redshift distribution modelling in Section 2.2 and of the luminosity-dependent modelling of IA in Section 2.4 above, we are in a position to explicitly introduce our proposed joint modelling framework.
@@ 2.6 Fisher forecasting
We explore the potential of our proposed joint-modelling parameter estimation framework via a Fisher forecasting analysis, chosen for its convenience and efficiency. In Fisher forecasting formalism, random variable $\boldsymbol{x}$ (an $N$-dimensional vector), has an associated probability density distribution $L(\boldsymbol{x};\boldsymbol{\theta})$ that depends on $M$ model parameters $\boldsymbol{\theta}=(\theta_{1},\theta_{2},...,\theta_{M})$. The central quantity is the Fisher matrix, defined as:
@@
$\mathscr{L}\equiv-\ln L$ (21)
The Fisher matrix can be computed without the need to locate the likelihood function. Instead, assuming that the probability distribution is a multivariate Gaussian, one can derive an analytical formula by examining the derivatives and their expectation values $\boldsymbol{\mu}\equiv\langle\boldsymbol{x}\rangle$ (Tegmark, 1997)
@@
with $\boldsymbol{C}\equiv\langle\boldsymbol{xx}^{t}\rangle-\langle\boldsymbol{\mu\mu}^{t}\rangle$ the covariance matrix and all quantities being evaluated at a set of fiducial values of the parameters $\boldsymbol{\theta}$. In the case in which the covariance matrix does not depend appreciably on the parameters $\boldsymbol{\theta}$, the above reduces to:
@@
which is the formulation of the Fisher matrix we use in this work. The Fisher matrix (referenced as Eq. (20)) and its inverse $\boldsymbol{F}^{-1}$ are the ideal matrices for error estimation in parameter measurements. According to the Cramér-Rao inequality, the minimum standard deviation achievable in measuring the $\alpha^{\mathrm{th}}$ parameter, with all others fixed, is $1/\sqrt{F_{\alpha\alpha}}$. However, if all parameters vary, the smallest possible error is $1/\sqrt{F_{\alpha\alpha}^{-1}}$ (Tegmark, 1997). Confidence contours, determined by the Fisher matrix, maintain constant probability in parameter space and are inherently ellipsoidal due to their Gaussian nature. These ellipsoids are described by the equation:
$\Delta\Phi^{\mathrm{T}}F\Delta\Phi=\beta$ (24)
where $\Delta\Phi=\Phi-\tilde{\Phi}$ represents the deviation from the best-fit value, and $\beta$ is a constant influenced by the desired confidence level and the count of free parameters.
In our study, we acknowledge the inherent instabilities in Fisher forecasting analyses due to the reliance on numerical derivatives, as highlighted in (Bhandari et al., 2021). In order to ensure the numerical stability, we employ the "stem method", as detailed in (Camera et al., 2016), which has been effectively applied in Euclid forecasts (Blanchard et al., 2020; Bonici et al., 2023). A detailed explanation of the implementation of this method into our framework is given in Sec. 3 on Fisher forecasting infrastructure.
@@ 3 Forecasting set-up
We now summarize the analysis choices and major framework building blocks which underpin the results of the remainder of this section. The forecasts we conduct are for the cosmic shear probe for LSST Year 1 and Year 10 (Y1 and Y10 hereafter). For a comprehensive overview of the relevant parameters and their values, please refer to Table 2.
Table 2: Modelling choices used in this analysis. For parameters which are varied in the analysis, the values given are fiducial. If not explicitly indicated, the central values are kept the same for both forecasting years. Values in parentheses are the $1\sigma$ width of the Gaussian prior imposed on this parameter in the analysis unless otherwise stated. No value in parentheses indicates a flat prior. Luminosity function parameters for all galaxies are the same as indicated in Table 1.
@@
Luminosity function model. The core objective of this work is to simultaneously model the redshift distribution and intrinsic alignments using a single luminosity function, aiming to produce LSST forecasts for Years 1 and 10. This entails finding fiducial luminosity function parameters which achieve congruence in both the redshift distributions and the IA amplitude with LSST expectations. To this end, we have calibrated the fiducial Schechter function parameters to achieve good agreement with the expected LSST redshift distribution and IA amplitude as defined in Mandelbaum et al. (2018). We use the parameters of the GAMA survey as a starting point representing a viable physical luminosity function, albeit from a survey with a very different selection. The direct application of GAMA parameters, unsurprisingly, revealed inadequacies in matching the LSST’s redshift distributions and IA amplitude using our framework (illustrated in Fig. 10), necessitating a thorough adjustment of these parameters. Consequently, the luminosity function parameters for both red and all (red+blue) galaxy populations were modified from the GAMA starting point. Through this process, we derive a fiducial luminosity function and consequently a red galaxy fraction that, while producing good agreement with the entire redshift distribution and redshift-dependent IA amplitude expected for LSST, diverge from a strict physical interpretation of ‘red’ galaxies and ‘all’ galaxies.
To be more concrete: our IA model, being a tidal alignment model, is best suited for red galaxies. However, in this work, we use a single set of luminosity function parameters for the ‘red’ luminosity function of Eqs. 10 and 11 as well as for the luminosity function $\Phi$ of Eq. 6 used in modelling the overall redshift distribution; these are the parameters for which we seek fiducial values which achieve agreement with LSST expectations and which we vary in our Fisher analysis below. For the parameters of the ‘all’ luminosity function of Eqs. 10 and 11, we use the ‘all’ parameters of GAMA directly and do not vary these in our analysis. This is technically unphysical, but produces fiducial redshift distribution, IA amplitude, and red- fraction which agree to an acceptable degree with our expectations for LSST. These modifications blur the distinction of what is categorically ‘red’ and ‘red fraction,’ rendering them less physically definitive yet crucial for our aims. These modifications were adopted as this work serves as a proof of concept, emphasizing that the project’s validity hinges on its ability to replicate realistic IA amplitudes and redshift distributions, aligning closely with the specifications detailed in the SRD, rather than on maintaining strict physical accuracies in galaxy classification. The refined values of these parameters, tailored to meet the forecasting scenarios, are detailed in Table 2. We acknowledge that this approach, while effective for our immediate goals, undoubtedly necessitates further investigation. A fully consistent treatment would accommodate sub-selections of the entire galaxy sample as being subject to different alignment effects; we leave the development of this full treatment for future work beyond this proof of concept.
@@
Redshift distribution and tomographic binning. We allow for the possibility that the galaxies in our cosmic shear analyses (source galaxies, ‘sources’ hereafter) span a redshift range of $0.01\leq z\leq 3.51$. In both forecast years, the redshift range step size is 0.01. The total redshift distribution of source galaxies as given in the LSST SRD is parameterised as:
$\left(\frac{\operatorname{d}\!{N}}{\operatorname{d}\!{z}}\right)\propto{\left(\frac{z}{z_{0}}\right)}^{2}\exp{\left(-\frac{z}{z_{0}}\right)^{\alpha}}$ (26)
with $(z_{0},\alpha)=(0.13,0.78)$ for Y1 and $(z_{0},\alpha)=(0.11,0.68)$ for Y10 (Mandelbaum et al., 2018).
@@
In our joint-modelling framework, we instead model the redshift distribution via the convolution method by van Daalen & White 2018, described in Sec. 2.2. We thus need to ensure that our fiducial luminosity function parameters reproduce a realistic source galaxy redshift distribution, considering the survey i-band limiting magnitudes of 24.1 and 25.3 for forecasting years 1 and 10, respectively. Our procedure for doing so is as follows: we start from the fixed survey volume ($18,000\;\mathrm{deg^{2}}$) and a luminosity function in a Schechter form. We initially take a set of LF parameters for $r$-band red galaxies from the GAMA survey (Loveday et al., 2012). From there, we use the insights gained from exploring the LF parameter space (explained in detail in Section 4.1) to find a set of fiducial LF parameters which produce redshift distributions which match the LSST Y1 and Y10 distributions as given in Mandelbaum et al. 2018. Specifically, we match the mean, mode, and median redshifts to the greatest extent possible. Apart from that, we inspect the corresponding cumulative distribution functions (CDFs) to ensure that the overall distribution shape aligns well with our expectations and the parameters of the study. To achieve this, we adjusted the density parameter $P$, rate of change of the pivot magnitude (luminosity) $Q$, as well as the faint end slope $\alpha$ from their initial GAMA survey values. The precise values and adjustments for these parameters can be found in Table 2. The redshift distributions we generated are consistent with the LSST SRD ones to a high degree, as can be seen in the upper panels of Fig. 2.
For both Y1 and Y10 forecasts, we follow the LSST binning routine (Mandelbaum et al., 2018) and segment the redshift distributions into five equipopulated tomographic bins. We operate under the premise that every PZ model can be represented by specifying a function for the distribution of photometric redshifts, given the true redshift $p(z_{\mathrm{PZ}}|z)$. Furthermore, we assume that this true redshift conforms to a Gaussian distribution at each redshift point (Ma et al., 2006):
$p(z_{\mathrm{PZ}}|z)=\frac{1}{\sqrt{2\pi}\sigma_{z}}\exp\left[-\frac{(z-z_{\mathrm{PZ}}-z_{\mathrm{bias}})^{2}}{\sigma_{z}^{2}}\right]$ (27)
with $z_{\mathrm{bias}}$ and $\sigma_{z}$ acting as the probability distribution parameters. The redshift distribution of tomographic bin $i$ is
$N_{i}(z)=0.5N(z)[\operatorname{erf}{(x_{i+1)}-\operatorname{erf}(x_{i})}]$ (28)
with
$x_{i}\equiv\frac{z^{(i)}_{\mathrm{PZ}}-z+z_{\mathrm{bias}}}{\sqrt{2}\sigma_{z}}$ (29)
@@
and $\operatorname{erf}{(x)}$ the error function. Each tomographic bin is then described by distinct uncertainty parameters. In the case of LSST, bias on the redshift $z_{\mathrm{bias}}$ is set to zero while scatter is a linear function of redshift
$\sigma_{z}=0.05(1+z)$ (30)
@@
for both forecast years. This is visualized in the bottom panels of Fig. 2, using green color for Y1 and blue for Y10. We also over-plot the LSST SRD bins (in dashed lines) for completeness and easy comparison.
@@
Figure 2: Source sample redshift distributions and tomographic bins used in this work. Top panels: Normalized source galaxies redshift distributions for LSST Y1 (left) and Y10 (right). Redshift distributions from Mandelbaum et al. 2018 (in dashed lines, labeled as "SRD", "Science Requirement Document"), against which we compare our fiducial distributions computed using the convolutional modelling method, are represented in solid lines in both plots. Corresponding redshift distributions used in our analyses (labeled as "JMAS","Joint Modelling of Astrophysics Systematics") are represented in green (Y1) and blue (Y10). The success of our implementation is corroborated by comparing the respective cumulative distribution functions (CDFs), indicated by dotted lines for CDFs from Mandelbaum et al. 2018 and solid green and blue lines for the method of this work, for Y1 and Y10, respectively. We also highlight the difference in the redshift distributions between Y1 and Y10 scenarios by including Y10 distribution (blue solid line) in the Y1 panel (top left) and Y1 distribution (green solid line) in the Y10 panel (top right). This difference plays a major role in our approach when interpreting the Fisher forecasting results, as highlighted in Sec. 4.4 Bottom panel: SRD (dashed lines) and JMAS tomographic bins for LSST Y1 and Y10 (green and blue shaded regions, respectively).
@@
Intrinsic alignment model. For the intrinsic alignments, our model incorporated a luminosity and redshift-dependent amplitude, as detailed in Eq. 8. The fiducial intrinsic alignment parameters — $A_{0}$, $\beta$, $\eta_{z}^{\mathrm{low}}$, and $\eta_{z}^{\mathrm{high}}$ — employed in our analysis are the fiducial values of Mandelbaum et al. 2018, adopted from Hirata & Seljak 2004a, Bridle & King 2007, and Krause et al. 2015, and can be found in Table 2. The average IA amplitude was integrated across both the magnitude range and redshifts, as explained in Krause et al. (2015). Furthermore, during our calculations, we accounted for the red fraction of galaxies (inset in Fig. 3) and the evolution of the IA amplitude across different redshifts was achieved by applying both low and high redshift scaling. This fiducial redshift evolution of the IA amplitudes can be viewed in Fig. 3: the solid green line represents the amplitude for Y1 while Y10 is represented in blue. The intrinsic alignment amplitude in both survey scenarios exhibits a sharp increase up to a redshift of approximately 1.5, beyond which it demonstrates a more gradual decline. Notably, the overall amplitude observed in the Y1 sample surpasses that projected for the year 10 forecasting scenario. For completeness and easier comparison, we included the redshift evolution of the IA signal for the corresponding SRD cases in Fig. 3.

### v3

@@ 2.3 Intrinsic alignment of galaxies
The framework proposed in this work can be applied to any intrinsic alignment model which incorporates luminosity dependence. Here, we choose the non-linear linear alignment (NLA) model (Bridle & King, 2007). We adopt a redshift- and luminosity-dependent parametrization for the NLA IA amplitude $A$, following Joachimi et al. (2011) and Krause et al. (2016):
$A(L,z)$ $=A_{0}\frac{C_{1}\rho_{m,0}}{D(z)}\left(\frac{L}{L_{0}}\right)^{\beta}\left(\frac{1+z}{1+z_{0}^{\mathrm{low}}}\right)^{\eta_{z}^{\mathrm{low}}}$
×$\times\left[\Theta(z_{0}^{\mathrm{high}}-z)+\Theta(z-z_{0}^{\mathrm{high}})\left(\frac{1+z}{1+z_{0}^{\mathrm{high}}}\right)^{\eta_{z}^{\mathrm{high}}}\right].$ (8)
$A_{0}$, $\beta$, $\eta_{z}^{\mathrm{low}}$, and $\eta_{z}^{\mathrm{high}}$ are free parameters of the model to be constrained, respectively capturing information about the amplitude of the IA signal, its scaling with luminosity, and its scaling with redshift in lower and higher $z$ regimes. $z_{0}^{\mathrm{low}}$ is a fixed pivot redshift of the low-redshift power law, while $z_{0}^{\mathrm{high}}$ defines the redshift at which the higher-redshift behavior (defined by $\eta_{z}^{\mathrm{high}}$) becomes important; similarly $L_{0}$ is a fixed pivot luminosity. The fiducial values for these parameters used in our forecasting analysis are provided in Table 2 (with $L_{0}$ given in terms of equivalent absolute magnitude, $M_{\rm piv}$). $C_{1}$ is a normalization constant (fixed to $C_{1}\rho_{\mathrm{cr}}\approx 0.0134$ via SuperCOSMOS observations (Hirata & Seljak, 2004), with $\rho_{\mathrm{cr}}$ the critical density), and $\rho_{\mathrm{m},0}$ is the matter density at $z=0$ (today). $D(z)$ is the linear growth factor, and $\Theta$ is a Heaviside step function.
We follow Krause et al. (2016) to obtain the IA amplitude as a function of redshift and survey limiting magnitude, by integrating over the luminosity function for the population of red galaxies:
$A(m_{\mathrm{lim}},z)=\biggl{\langle}A(L,z)\biggr{\rangle}_{{\phi_{\mathrm{red}}}}{f_{\rm red}(z)}.$ (9)
Here, $m_{\mathrm{lim}}$ is the survey limiting magnitude and $f_{\rm red}(z)$ is the fraction of red galaxies as a function of redshift. Explicitly, the angular brackets in Eq. 9 are evaluated as:
@@
$\biggl{\langle}A(L,$ $z)\biggr{\rangle}_{\phi_{\mathrm{red}}}{=A(L_{0},z)\left\langle\left(\frac{L}{L_{0}}\right)^{\beta}\right\rangle_{\phi_{\mathrm{red}}}\left(\frac{1+z}{1+z_{0}^{\mathrm{low}}}\right)^{\eta_{z}^{\mathrm{low}}}}$
×${\times\left[\Theta(z_{0}^{\mathrm{high}}-z)+\Theta(z-z_{0}^{\mathrm{high}})\left(\frac{1+z}{1+z_{0}^{\mathrm{high}}}\right)^{\eta_{z}^{\mathrm{high}}}\right]}$
${=A(L_{0},z)\frac{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\left(\frac{L}{L_{0}}\right)^{\beta}\phi_{\mathrm{red}}(L,z)}{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\phi_{\mathrm{red}}(L,z)}\left(\frac{1+z}{1+z_{0}^{\mathrm{low}}}\right)^{\eta_{z}^{\mathrm{low}}}}$
×${\times\left[\Theta(z_{0}^{\mathrm{high}}-z)+\Theta(z-z_{0}^{\mathrm{high}})\left(\frac{1+z}{1+z_{0}^{\mathrm{high}}}\right)^{\eta_{z}^{\mathrm{high}}}\right]}.$ (10)
Finally, the fraction of red galaxies, is obtained via:
$f_{\mathrm{red}}{(z)}=\frac{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\phi_{\mathrm{red}}(L,z)}{\int^{\infty}_{L(m_{\mathrm{lim}},z)}\operatorname{d}\!{L}\phi_{\mathrm{all}}(L,z)}$ (11)
with $\phi_{\mathrm{red}}$ and $\phi_{\mathrm{all}}$ the luminosity functions of red and all galaxies in the sample.
@@
An important emphasis is that the IA amplitude is modelled as luminosity-dependent. Thus, when computing the cosmic shear spectrum from theory, it is necessary to assume a luminosity function parametrization. Previous forecasting work using this parametrization has fixed the luminosity function parameters to best-fit values measured from external data sets (Krause et al., 2016; Mandelbaum et al., 2018). However, it is not clear that this is an appropriate analysis choice; we will investigate this as part of the development of our analysis framework below in Sec. 3.
@@ 2.4 Cosmic shear
Each contribution is described in spherical harmonic space where $\ell$ is the angular frequency (a 2D wave vector perpendicular to the line of sight) while indices $(i,j)$ represent tomographic redshift bins where the shear correlations are computed (for a review of key intrinsic alignment quantities and notation, see Lamman et al. (2024)).
In this work, we will take our fiducial IA model to be the non-linear alignment (NLA) model (Bridle & King, 2007; Krause et al., 2016). We stress that although it is likely that a more complex model of IA may be required for LSST analysis (e.g. the TATT model from Blazek et al. 2019), the NLA model provides a suitable means of incorporating luminosity-dependent IA modelling and is appropriate in the context of a first exploration of our joint modelling methodology. For the case of the non-linear alignment model of intrinsic alignments, each term in the observed cosmic shear power spectrum is given by:
@@
where in the linear alignment model, $P_{\epsilon\epsilon}$ and $P_{\delta\epsilon}$ are given by the linear matter power spectrum, whilst the non-linear alignment model uses the non-linear matter power spectrum. Both linear and non-linear matter power spectra are expressed in the Fourier space, $k$. $A(z)$ is the redshift-dependent intrinsic alignment amplitude that also depends on the galaxy properties and we use the amplitude parametrization presented in Sec. 2.3.
@@ 2.5 Joint modelling via the luminosity function
Given the introduction of the method of modelling the redshift distribution of galaxies via a volume selection and luminosity function in Section 2.2 and of the luminosity-dependent modelling of IA in Section 2.4 above, we are in a position to explicitly introduce our proposed joint modelling framework.
@@ 2.6 Fisher forecasting
We explore the potential of our proposed joint-modelling parameter estimation framework via a Fisher forecasting analysis, chosen for its convenience and efficiency. In the Fisher forecasting formalism, random variable $\boldsymbol{x}$ (an $N$-dimensional vector) has an associated probability density distribution $L(\boldsymbol{x};\boldsymbol{\theta})$ that depends on $M$ model parameters $\boldsymbol{\theta}=(\theta_{1},\theta_{2},...,\theta_{M})$. The central quantity is the Fisher matrix, defined as:
@@
$\mathscr{L}\equiv-\ln L.$ (21)
Assuming that the probability distribution is a multivariate Gaussian, one can derive an analytical formula by examining the derivatives and their expectation values $\boldsymbol{\mu}\equiv\langle\boldsymbol{x}\rangle$ (Tegmark, 1997)
@@
with $\boldsymbol{C}\equiv\langle\boldsymbol{xx}^{t}\rangle-\boldsymbol{\mu\mu}^{t}$ the covariance matrix and all quantities being evaluated at a set of fiducial values of the parameters $\boldsymbol{\theta}$. In the case in which the covariance matrix does not depend appreciably on the parameters $\boldsymbol{\theta}$, the above reduces to:
@@
which is the formulation of the Fisher matrix we use in this work. According to the multivariate Cramér-Rao inequality (Rao, 1992), the inverse of the Fisher matrix is the lower bound on the covariance matrix of the cosmological parameters (under the assumption that our analysis set-up allows for an unbiased estimate of the parameters from data):
$\text{cov}(\hat{\boldsymbol{\theta}})\geq\boldsymbol{F}^{-1}\,.$ (24)
Here $\hat{\boldsymbol{\theta}}$ denotes the vector of estimated parameter values and $\boldsymbol{F}^{-1}$ denotes the inverse of the Fisher information matrix.
We note that in this work we incorporate prior information in our Fisher forecasts (with the parameters of these Gaussian priors given below in Table 2). We follow the methodology of e.g. Bassett et al. (2011) in assuming such prior information is independent from the information offered in the present analysis, such that we can simply add Fisher matrices representing this prior information to our Fisher matrix before inverting to obtain estimated (lower bounds on the) parameter covariance.
In our study, we acknowledge the inherent instabilities in Fisher forecasting analyses due to the reliance on numerical derivatives, as highlighted in Bhandari et al. (2021). In order to ensure the numerical stability, we employ the "stem method", as detailed in Camera et al. (2016), which has been effectively applied in Euclid forecasts (Blanchard et al., 2020; Bonici et al., 2023). A detailed explanation of the implementation of this method into our framework is given in Sec. 3 on Fisher forecasting infrastructure.
@@ 3 Forecasting set-up
We now summarize the analysis choices and major framework building blocks which underpin our results. The forecasts we conduct are for the cosmic shear probe for LSST Year 1 and Year 10 (Y1 and Y10 hereafter). For a comprehensive overview of the relevant parameters and their values, please refer to Table 2.
Table 2: Modelling choices used in this analysis. For parameters which are varied in the analysis, the values given are fiducial. If not explicitly indicated, the central values are kept the same for both forecasting years. Values in parentheses are the $1\sigma$ width of the Gaussian prior imposed on this parameter in the analysis unless otherwise stated. No value in parentheses indicates a flat prior. Luminosity function parameters for ‘all’ galaxies are the same as indicated in Table 1.
@@
Luminosity function model. We allow a single set of luminosity function parameters to describe the ‘red’ luminosity function of Eqs. 10 and 11 (used in modelling IA) and the total luminosity function of Eq. 6 (used in modelling the redshift distribution). We take this set of parameters as free parameters to be constrained within a cosmic shear analysis. For the parameters of the ‘all’ luminosity function of Eqs. 10 and 11 (used in modelling IA), we use the ‘all’ parameters of GAMA directly (given in Table 2) and do not vary these in our analysis. This set-up is appropriate for our objective of a simplified, proof-of-concept analysis in which the IA amplitude and redshift distribution modelling share a set of luminosity function parameters, while the red fraction (and hence the fractional impact of IA on cosmic shear) is kept to a physically expected level for LSST. A more complete and correct treatment would incorporate additionally a selection function for color as a function of luminosity and redshift, with additional free parameters to be constrained; we leave the explicit incorporation of such selection functions for future work.
Given this set-up, we require fiducial values of free parameters of our luminosity function which achieve both the redshift distributions and IA amplitude which agree with LSST expectations as defined in Mandelbaum et al. (2018). We follow Krause et al. (2016) in using parameter values as fit to the GAMA survey as a starting point (representing as they do a viable physical luminosity function, albeit from a survey with a very different selection). Given this considerable difference in survey selection (with GAMA having a much brighter limiting magnitude than that expected for LSST Y1 or Y10), it is not surprising that directly implementing these GAMA luminosity function parameters does not result in redshift distributions or IA amplitudes which match those expected for LSST (illustrated in Fig. 12). Given this, we iteratively adjust the luminosity function parameters to achieve a fiducial luminosity function which, within our modelling framework, produces good agreement with the redshift distribution and redshift-dependent IA amplitude expected for LSST. The refined values of these parameters, tailored to meet the forecasting scenarios, are detailed in Table 2.
@@
Redshift distribution and tomographic binning. We allow for the possibility that the galaxies in our cosmic shear analyses (source galaxies, ‘sources’ hereafter) span a redshift range of $0.01\leq z\leq 3.51$. In both forecast years, the redshift range step size is 0.01. The total redshift distribution of source galaxies as given in the LSST Science Requirements Document (SRD) is parametrized as:
$\frac{\operatorname{d}\!{N}}{\operatorname{d}\!{z}}\propto{\left(\frac{z}{z_{0}}\right)}^{2}\exp{\left[-\left(\frac{z}{z_{0}}\right)^{\alpha}\right]}$ (26)
@@
with $(z_{0},\alpha)=(0.13,0.78)$ for Y1 and $(z_{0},\alpha)=(0.11,0.68)$ for Y10 (Mandelbaum et al., 2018). In our joint-modelling framework, we instead model this redshift distribution via the method by van Daalen & White (2018), described in Sec. 2.2. We thus need to ensure that our fiducial luminosity function parameters reproduce a realistic underlying, unbinned source galaxy redshift distribution, considering the survey i-band limiting magnitudes of 24.1 and 25.3 for forecasting years 1 and 10, respectively. Our procedure for doing so is as follows: we start from a luminosity function in a Schechter form. We initially take a set of LF parameters for $r$-band red galaxies from the GAMA survey (Loveday et al., 2012). From there, we use the insights gained from exploring the LF parameter space (explained in detail in Section 4.1) to find a set of fiducial LF parameters which produce redshift distributions which match the LSST Y1 and Y10 distributions as given in Mandelbaum et al. (2018). Specifically, we match the mean, mode, and median redshifts to the greatest extent possible. Apart from that, we inspect the corresponding cumulative distribution functions (CDFs) to ensure that the overall distribution shape aligns well with our expectations and the parameters of the study. To achieve this, we adjusted the density parameter $P$, rate of change of the pivot magnitude (luminosity) $Q$, as well as the faint end slope $\alpha$ from their initial GAMA survey values. The precise values and adjustments for these parameters can be found in Table 2. The redshift distributions we generated are consistent with the LSST SRD ones to a high degree, as can be seen in the upper panels of Fig. 2.
For both Y1 and Y10 forecasts, we then follow Mandelbaum et al. (2018) to create five tomographic bins and to incorporate effects on the binning from photometric redshift uncertainty. We first segment the overall true redshift distribution into five equipopulated tomographic bins by applying top-hat window functions constructed on this basis, $T_{i}(z)$, for each bin $i$. We then take the simplifying assumption that the joint probability distribution of the true redshift with the photometric redshift in each tomographic bin $i$ conforms to a Gaussian distribution at each redshift point (see, for example, Ma et al. 2006):
$p^{i}(z_{\mathrm{PZ}}|z)=\frac{1}{\sqrt{2\pi}{\sigma_{z}^{i}}}\exp\left[-\frac{(z-z_{\mathrm{PZ}}-z_{\mathrm{bias}}^{i})^{2}}{2(\sigma_{z}^{i})^{2}}\right].$ (27)
with $z_{\mathrm{bias}}^{i}$ representing for redshift bin $i$ a shift in the mean of the Gaussian and away from $z=z_{PZ}$ and $\sigma_{z}^{i}$ representing the possibly-z-dependent variance term. Given this, we have the redshift distribution $\frac{dN}{dz}^{i}$ of tomographic bin $i$ to be:
[…]
@@
Each tomographic bin is then described in principle by distinct uncertainty parameters. Fiducially, bias on the redshift $z_{\mathrm{bias}}^{i}$ is set to zero in each bin $i$ while scatter in the LSST case is modelled for all bins as a linear function of redshift
$\sigma_{z}^{i}=0.05(1+z)$ (29)
@@
for both forecast years. Although it is conventional to vary one or both of these sets of parameters in cosmological analysis, we fix them here. This is because we instead allow to vary the luminosity parameters which produce to underlying redshift distribution. The resulting redshift distributions are visualized in the bottom panels of Fig. 2, using green color for Y1 and blue for Y10. We also over-plot the LSST SRD bins (in dashed lines) for completeness and easy comparison.
@@
Figure 2: Source sample redshift distributions and tomographic bins used in this work. Top panels: Normalized source galaxies redshift distributions for LSST Y1 (left) and Y10 (right). Redshift distributions from Mandelbaum et al. (2018) (in dashed lines, labeled as "SRD", "Science Requirement Document"), against which we compare our fiducial distributions computed using the modelling method described in Sec. 2.2, are represented in solid lines in both plots. Corresponding redshift distributions used in our analyses (labeled as "JMAS","Joint Modelling of Astrophysics Systematics") are represented in green (Y1) and blue (Y10). The success of our implementation is corroborated by comparing the respective cumulative distribution functions (CDFs), indicated by dotted lines for CDFs from Mandelbaum et al. (2018) and solid green and blue lines for the method of this work, for Y1 and Y10, respectively. We also highlight the difference in the redshift distributions between Y1 and Y10 scenarios by including Y10 distribution (blue solid line) in the Y1 panel (top left) and Y1 distribution (green solid line) in the Y10 panel (top right). This difference plays a major role in our approach when interpreting the Fisher forecasting results, as highlighted in Sec. 4.4 Bottom panel: SRD (dashed lines) and JMAS tomographic bins for LSST Y1 and Y10 (green and blue shaded regions, respectively).
@@
Intrinsic alignment model. For the intrinsic alignments, our model incorporated a luminosity and redshift-dependent amplitude, as detailed in Eq. 8. The fiducial intrinsic alignment parameters — $A_{0}$, $\beta$, $\eta_{z}^{\mathrm{low}}$, and $\eta_{z}^{\mathrm{high}}$ — employed in our analysis are the fiducial values of Mandelbaum et al. (2018) and can be found in Table 2. The fiducial redshift evolution of the IA amplitudes as well as the corresponding red fraction can be viewed in Fig. 3: the solid green line represents the amplitude for Y1 while Y10 is represented in blue. The intrinsic alignment amplitude in both survey scenarios exhibits a sharp increase up to a redshift of approximately 1.5, beyond which it demonstrates a more gradual decline. Notably, the overall amplitude observed in the Y1 sample surpasses that projected for the year 10 forecasting scenario. For completeness and easier comparison, we included the redshift evolution of the IA signal for the corresponding SRD cases in Fig. 3.